A string on a cello has a tension of 34.5N. The linear density is 0.05g / m . What is the speed of sound along the string ? 


1
Expert's answer
2021-01-24T16:32:09-0500

"v=\sqrt{\frac{T}{\rho}}=\sqrt{\frac{34.5}{0.05\cdot10^{-3}}}=831(m/s)"